OpenStudy (tvinsant):

Solve for b. -23ab + b = 17a

OpenStudy (deepolisnoob):

This situation involves factoring. Firstly, you must get all your factors to one side, so you subtract 17a from both sides to give you:\[-17a-23ab+b=0\] After that, you factor your equation

OpenStudy (anonymous):

Correct @Deepolisnoob

OpenStudy (tvinsant):

What is the answer to this question? a. it is impossible to solve for b. b. b=-17/23 c. b=40a d. b=17a/-23a+1

OpenStudy (deepolisnoob):

You are not looking for the factors of the prior equation; to get the value of b, you must isolate b. Subtract b from both sides to get b by itself, giving you: \[-17a-23ab=-b\] You want to get a positive b rather than the negative b you have, you you multiply everything by -1, giving you positive values to replace the negative ones.

OpenStudy (tvinsant):

Ok? But which one is the answer? sorry

OpenStudy (tvinsant):

Because it doesn't seem it fit any of the options.

OpenStudy (deepolisnoob):

b=17a+23ab should be the best fit...in your equation, are your a and b values squared?

OpenStudy (tvinsant):

no

OpenStudy (deepolisnoob):

I'm not sure what the answer could be; the simplest you can reach is b=a(17+23b). I don't agree with any of the answer choices provided.

OpenStudy (tvinsant):

Me neither.